how many grams of O2 must you have to produce 14.6 grams of H2O answwr


1
Expert's answer
2021-03-08T07:06:18-0500

The equation for the reaction

O2 + 2H2 "\to" 2H2O


Moles of water used "=\dfrac{mass}{RMM} = \dfrac{14.6g}{18.01528g/mol}=0.81042mol"


Mole ratio of O2 : H2O = 1:2


Moles of O2"=\dfrac{1}{2}x0.81042 = 0.4052mol"


Mass of O2 = moles x MM

= 0.4052mol x 31.998 =12.966g